OpenStudy (anonymous):

What is log (x+2) - log x = 2 ? Please Answer po . :DD

hartnn (hartnn):

you need to solve for 'x' right ? use the property that \(\log x-\log y = \log (x/y)\)

OpenStudy (anonymous):

yes po .. :D i already try that po ..

hartnn (hartnn):

so, which step are you stuck ?

OpenStudy (shubhamsrg):

po po po po po, po po po pooo! ^_^

OpenStudy (anonymous):

@shubhamsrg che! :D haha!

OpenStudy (anonymous):

i don't know po. hehe! :DD

hartnn (hartnn):

if you used that formula, you must be here, log [(x+2)/x ] =2 right ? now since log has base 10, use, \(\log a=b \implies a=10^b\) here a= [(x+2)/x ] ,b =2

OpenStudy (anonymous):

yeah .. i get it .. the naswer is 2/99 right??

hartnn (hartnn):

yes! :) good.

hartnn (hartnn):

any more doubts ?

OpenStudy (anonymous):

none po .. :d thank you po for helping me. GOD bless you always po . :D

hartnn (hartnn):

welcome ^_^ God bless you too :)

OpenStudy (anonymous):

naks! :D thank you again po. :D :D SAYOUNARA! XIE XIE. :D

hartnn (hartnn):

i know SAYOUNARA means bye :D good bye :)

hartnn (hartnn):

or am i incorrect ? O.o

OpenStudy (anonymous):

:D correct! :D clap! clap! clap! :D very good! :D haha!

hartnn (hartnn):

thank you thank you :D :)

OpenStudy (anonymous):

:D welcome welcome welcome po . :D haha! why are you so good po in math?

hartnn (hartnn):

because when i was studying maths, i did lots and lots of practice! thats the only secret for learning math....

OpenStudy (anonymous):

wow! :D hehe .. :D "MAGAYA NGA!" :D did you know that word po?

hartnn (hartnn):

no don't know.... but i can always use google translate ;)

OpenStudy (anonymous):

ahh .. hehe! it means "i will imitate that po" :D what is your nationality po?

hartnn (hartnn):

ohh..ok i am INDIAN .

OpenStudy (anonymous):

:D wow! :D nice.. okay po. :D hehe.. nice to meet you sir ? what is your real name po by the way?

hartnn (hartnn):

nice to meet you too innie :) my real name is Hardik :)

OpenStudy (anonymous):

:D ahh! hehe .. how did you know again po my nickname? hm .. i didn't get it last time when im asking you that po but i'm replying "YES" even i don't understand po . .hehe! :D